# Large Sets of Asymptotically Optimal and Near-Optimal Quasi-Complementary

### v1

@@ Abstract
Perfect complementary sequence sets (PCSSs) are widely used in multi-carrier code-division multiple-access (MC-CDMA) communication system. However, the set size of a PCSS is upper bounded by the number of row sequences of each two-dimensional matrix in PCSS. Then quasi-complementary sequence set (QCSS) was proposed to support more users in MC-CDMA communications. For practical applications, it is desirable to construct an $(M,K,N,\vartheta_{max})$-QCSS with $M$ as large as possible and $\vartheta_{max}$ as small as possible, where $M$ is the number of matrices with $K$ rows and $N$ columns in the set and $\vartheta_{max}$ denotes its periodic tolerance. There exists a tradoff among these parameters and constructing QCSSs achieving or nearly achieving the known correlation lower bound has been an interesting research topic. Up to now, only a few constructions of asymptotically optimal or near-optimal periodic QCSSs were reported in the literature. In this paper, we construct five families of asymptotically optimal or near-optimal periodic QCSSs with large set sizes and low periodic tolerances. These families of QCSSs have set size $\Theta(q^{2})$ or $\Theta(q^{3})$ and flock size $\Theta(q)$, where $q$ is a power of a prime. To the best of our knowledge, only three known families of periodic QCSSs with set size $\Theta(q^{2})$ and flock size $\Theta(q)$ were constructed and all other known periodic QCSSs have set sizes much smaller than $\Theta(q^{2})$. Our new constructed periodic QCSSs with set size $\Theta(q^{2})$ and flock size $\Theta(q)$ have better parameters than known ones. They have larger set sizes or lower periodic tolerances. The periodic QCSSs with set size $\Theta(q^{3})$ and flock size $\Theta(q)$ constructed in this paper have the largest set size among all known families of asymptotically optimal or near-optimal periodic QCSSs.
@@
Quasi-complementary sequence set, CDMA communication, correlation lower bound
@@ I Introduction
A perfect complementary sequence set (PCSS), which is also called mutually orthogonal complementary sequence set, consists of a number of two-dimensional matrices. The nontrivial auto-correlation and cross-correlation sums of the row sequences of the matrices are zero for any nonzero time-shift [21]. PCSSs have been found many nice applications including peak-to-average power ration reduction [4], inter-symbol interference channel estimation [20], radar waveform design [18] and multi-carrier code-division multiple-access (MC-CDMA) communication system [2]. The application of PCSSs has been viewed as a very promising technology for the next generation CDMA [3]. In MC-CDMA communication system, each user is assigned a two-dimensional matrix whose all row sequences are transmitted over separate sub-carrier channels simultaneously [2]. Then the number of the row sequences of a matrix in PCSS and that of sub-carrier channels are equal. However, the set size of a PCSS can not be larger than the number of row sequences of each two-dimensional matrix in it by the correlation lower bound. In order to support more users in MC-CDMA communication systems, some researchers proposed two new kinds of complementary sequence sets which are zero-correlation zone complementary sequence sets (ZCZ-CSSs) [5, 8] and low-correlation zone complementary sequence sets (LCZ-CSS) [13]. An LCZ-CSS (resp. ZCZ-CSS) has low (resp. zero) correlation magnitudes within a zone around the zero-shift. In recent years, many researches have been done on constructing LCZ-CSSs and ZCZ-CSSs.
@@
In 2013, Liu et al. proposed another type of complementary sequence set called quasi-complementary sequence set (QCSS) which could have larger set size than that of PCSS [14]. In the literature, there are two kinds of QCSSs including periodic QCSSs and aperiodic QCSSs [14, 15]. In this paper, we mainly study periodic QCSSs. A periodic $(M,K,N,\vartheta_{max})$-QCSS over an alphabet $A$ is a set consisting of $M$ two-dimensional $K\times N$ matrices such that the maximum value $\vartheta_{max}$ of nontrivial periodic correlation sums of the row sequences is a small positive number. Each row of the two-dimensional matrices is called a constituent sequence of length $N$ and $K$ denotes the flock size. The maximum non-trivial periodic correlation sum $\vartheta_{max}$ measures the maximum possible amount of multipath interference and multiuser interference. For the application of a periodic QCSSs in MC-CDMA communications, it is required that its set size $M$ is as large as possible and $\vartheta_{max}$ is as small as possible. However, there exists a tradeoff among the parameters $M,K,N$ and $\vartheta_{max}$. In [14], Liu et al. presented a lower bound for the maximum periodic correlation sum of a periodic $(M,K,N,\vartheta_{max})$-QCSS as follows:
$\vartheta_{max}\geq\vartheta_{opt}=KN\sqrt{\frac{M/K-1}{MN-1}}.$ (1)
It is easy to verify that $M\leq K$, which is the upper bound of the size of a periodic PCSS, provided that $\vartheta_{max}=0$. The tightness factor of the above correlation lower bound for periodic QCSS is defined by $\rho=\vartheta_{max}/\vartheta_{opt}$. Note that $\rho\geq 1$. A QCSS is said to be optimal if $\rho=1$ and near-optimal if $1<\rho\leq 2$ [14].
@@
Constructing optimal or near-optimal periodic QCSSs has been an interesting research topic in recent years. However, up to now, no infinite family of optimal QCSSs achieving the lower bound in (1) was reported. Only a few families of asymptotically optimal or near-optimal QCSSs were constructed. In [14], Liu et al. constructed the first family of asymptotically optimal and the first family of asymptotically near-optimal periodic QCSSs from the Single difference sets and the existing optimal quaternary sequence sets. Using a similar idea, Li et al. constructed a family of asymptotically optimal periodic QCSSs from almost difference sets in [12]. In [9], Li et al. generalized the constructions in [14] and derived new asymptotically optimal periodic QCSSs. In [10, 11], Li et al. presented several constructions of asymptotically optimal QCSSs from additive and multiplicative characters over finite fields. Later, Luo et al. gave three new constructions of asymptotically optimal QCSSs with samll alphabet sizes by additive characters over finite fields [16]. Recently, Xiao et al. used additive characters of finite fields to construct two more constructions of asymptotically optimal periodic QCSSs [22]. The parameters of these known families of periodic QCSSs are summarized in Table I. It is shown that only three known families of periodic QCSSs with set size $\Theta(q^{2})$ and flock size $\Theta(q)$ were constructed and all other known periodic QCSSs have set sizes much smaller than $\Theta(q^{2})$, where $q$ is a power of a prime.
The purpose of this paper is to use new approaches to construct five new families of asymptotically optimal or near optimal QCSSs with better parameters. The main contents of this paper are as follows:
  1. 1.
@@
    Using a family of quadratic polynomials over finite fields, we construct a family of $(q^{2},q,q-1,q)$-QCSSs with alphabet size $p$ and a family of $(2^{2n},2^{n},2^{2n}-1,2^{n})$-QCSSs with alphabet size $2$, where $q=p^{n}$ with $p$ an odd prime and $n$ a positive integer. These two families of QCSSs are asymptotically optimal with respect to the correlation lower bound and have larger set sizes than those in [10, 22] by Table I.
  2. 2.
@@
    A family of $(2^{3n},2^{n},2^{n}-1,2^{n+1})$-QCSSs with alphabet size $2$ is constructed from a special quadrinomial over finite fields with characteristic 2. A family of $(p^{3n},p^{n},p^{n}-1,\leq 2p^{n})$-QCSSs with alphabet size $p$ is also constructed from a cubic quadrinomial over finite fields with characteristic $p$ for an odd prime $p$. These two families of QCSSs are asymptotically near-optimal with respect to the correlation lower bound. Note that the set size of each family of QCSSs is the cube of the flock size. In this sense, these two families of QCSSs have the largest set sizes among all known families of asymptotically optimal or near-optimal periodic QCSSs.
  3. 3.
@@
TABLE I: The parameter of known periodic QCSSs.
Set size $M$ Flock size $K$ Length $N$ $\vartheta_{max}$ Alphabet Size Constraint References
@@
$p^{n}$ $\frac{p^{n}-1}{2}$ $p^{n}-1$ $\frac{p^{n}+1}{2}$ $p$ $p$ is an odd prime, $p^{n}>3$ [16]
$p^{n}$ $\frac{p^{n}-p^{n-1}}{2}$ $p^{n}-1$ $\frac{p^{n}+p^{n-1}}{2}$ $p$ $p$ is an odd prime, $p^{n}>3$ [16]
$p^{n}$ $u$ $p^{n}-1$ $v$ $p$ $p$ is an odd prime, $n>1$ is odd [16]
$p^{n}(p^{n}-1)$ $p^{n}$ $p^{n}-1$ $p^{n}$ $p$ $p$ is a prime, $n>1$ [22]
$p^{n}(p^{n}-1)$ $p^{n}-1$ $p^{n}-1$ $p^{n}+1$ $p$ $p$ is a prime, $n>1$ [22]
@@
$p^{3n}$ $p^{n}$ $p^{n}-1$ $2p^{n}$ $p$ $p$ is an odd prime, $n\geq 1$ Theorem III.4
$p^{n}(p^{n}-1)$ $p^{n}-1$ $p^{n}-1$ $p^{n}$ $p(p^{n}-1)$ $p$ is a prime Theorem III.5
  • $u=\frac{p^{n}-p^{n-1}+(-1)^{(p-1)(n+3)/4}(p-1)p^{(n-1)/2}}{2}$,$v=\frac{p^{n}+p^{n-1}-(-1)^{(p-1)(n+3)/4}(p-1)p^{(n-1)/2}}{2}$,$w=(1+2^{n})\sqrt{2^{n}+2^{n-2}-2^{n/2}}$
@@ II Preliminaries
In this section, we recall some basic definitions of QCSSs and results on characters over finite fields, Gaussian sums, character sums, and the numbers of zeros of some equations over finite fields.
@@ II-A Periodic Quasi-Complementary Sequence Set
Let =$\mathbb{C}=\{\mathbf{C}^{0},\mathbf{C}^{1},\cdots,\mathbf{C}^{M-1}\}$ be a set of $M$ complementary sequences such that each $\mathbf{C}^{m}$ is a two-dimensional matrix with size $K\times N$ given by
@@
$R_{{{\textbf{C}}^{{m}_{1}}},{{\textbf{C}}^{{m}_{2}}}}(\tau)=\sum_{k=0}^{K-1}R_{{\mathbf{c}_{k}^{{m}_{1}}},{\mathbf{c}_{k}^{{m}_{2}}}}(\tau),~{}0\leq\tau<N,$
@@
$\vartheta_{a}=max\{|R_{{{\textbf{C}}^{{m}}},{{\textbf{C}}^{{m}}}}(\tau)|:0\leq m<M,0<\tau<N\},$
$\vartheta_{c}=max\{|R_{{{\textbf{C}}^{{m}_{1}}},{{\textbf{C}}^{{m}_{2}}}}(\tau)|:0\leq m_{1}\neq m_{2}<M,0\leq\tau<N\}.$
@@
$\vartheta_{max}=max\{\vartheta_{c},\vartheta_{a}\}.$
It is an important performance measure of the complementary sequence set $\mathbb{C}$ in practice applications. If $\vartheta_{max}>0$, then $\mathbb{C}$ is called a periodic $(M,K,N,\vartheta_{max})$ quasi-complementary sequence set (QCSS). Otherwise, $\mathbb{C}$ is called a periodic perfect complementary sequence set (PCSS). Particularly, a PCSS reduces to a matrix consisting of two row sequences provided that $M=1$ and $K=2$. This matrix is called a Golay complementary pair.
@@ II-B Additive and Multiplicative Characters of finite fields
for each $j=0,1,\cdots,q-2$ defines a multiplicative character. In particular, $\varphi_{0}$ is called the canonical multiplicative character. For odd $q$, if $j=\frac{q-1}{2}$, then $\eta:=\varphi_{\frac{q-1}{2}}$ is called the quadratic multiplicative character of ${\mathbb{F}}_{q}$. The complex conjugate $\overline{\varphi}$ of a multiplicative character $\varphi$ is defined by $\overline{\varphi}(x)=\overline{\varphi(x)}=\varphi(x^{-1})$ for $x\in{\mathbb{F}}_{q}^{*}$.
From [17], the orthogonal relation of multiplicative characters is given by
$\sum_{x\in{\mathbb{F}}_{q}^{*}}\psi_{j}(x)=\left\{\begin{array}[]{ll}q-1&\mbox{if $j=0$},\\ 0&\mbox{if $1\leq j\leq q-2$}.\end{array}\right.$
@@ II-C Gaussian sums and character sums over finite fields
In the following, we present some known results on the numbers of zeros of some equations over finite fields.
Lemma II.4

### v2

# Large Sets of Quasi-Complementary Sequences From Polynomials over Finite Fields and

@@ Abstract
Perfect complementary sequence sets (PCSSs) are widely used in multi-carrier code-division multiple-access (MC-CDMA) communication systems. However, the set size of a PCSS is upper bounded by the number of row sequences of each two-dimensional matrix in the PCSS. Then quasi-complementary sequence sets (QCSSs) were proposed to support more users in MC-CDMA communications. For practical applications, it is desirable to construct an $(M,K,N,\vartheta_{\max})$-QCSS with $M$ as large as possible and $\vartheta_{max}$ as small as possible, where $M$ is the number of matrices with $K$ rows and $N$ columns in the set and $\vartheta_{\max}$ denotes its periodic tolerance. There exists a tradeoff among these parameters. Constructing QCSSs achieving or nearly achieving the known correlation lower bound has been an interesting research topic. Up to now, only a few constructions of asymptotically optimal or near-optimal periodic QCSSs have been reported in the literature. In this paper, based on polynomials over finite fields and Gaussian sums, we construct five new families of asymptotically optimal or near-optimal periodic QCSSs with large set sizes and low periodic tolerances. These families of QCSSs have set size $\Theta(K^{2})$ or $\Theta(K^{3})$ and flock size $K$. To the best of our knowledge, only a small amount of known families of periodic QCSSs with set size $\Theta(K^{2})$ have been constructed and most of other known periodic QCSSs have set sizes much smaller than $\Theta(K^{2})$. Our new constructed periodic QCSSs with set size $\Theta(K^{2})$ and flock size $K$ have the best parameters among all known ones. They have larger set sizes or lower periodic tolerances. The periodic QCSSs with set size $\Theta(K^{3})$ and flock size $K$ constructed in this paper have the largest set size among all known families of asymptotically optimal or near-optimal periodic QCSSs.
@@
Quasi-complementary sequence sets, polynomials over finite fields, Gaussian sums
@@ I Introduction
A perfect complementary sequence set (PCSS), which is also called mutually orthogonal complementary sequence set, consists of a number of two-dimensional matrices. The nontrivial auto-correlation and cross-correlation sums of the row sequences of the matrices are zero for any nonzero time-shift [21]. PCSSs have found many nice applications including peak-to-average power ration reduction [4], inter-symbol interference channel estimation [20], radar waveform design [18] and multi-carrier code-division multiple-access (MC-CDMA) communication systems [2]. The application of PCSSs has been viewed as a very promising technology for the next generation CDMA [3]. In an MC-CDMA communication system, each user is assigned a two-dimensional matrix and all of its row sequences are transmitted over separate sub-carrier channels simultaneously [2]. Then the number of the row sequences of a matrix in the PCSS and that of sub-carrier channels are equal. However, the set size of a PCSS can not be larger than the number of row sequences of each two-dimensional matrix in it by the correlation lower bound. In order to support more users in MC-CDMA communication systems, some researchers proposed two new kinds of complementary sequence sets which are zero-correlation zone complementary sequence sets (ZCZ-CSSs) [5, 8] and low-correlation zone complementary sequence sets (LCZ-CSS) [13]. An LCZ-CSS (resp. ZCZ-CSS) has low (resp. zero) correlation magnitudes within a zone around the zero-shift. In recent years, many researches have been done on constructing LCZ-CSSs and ZCZ-CSSs.
@@
In 2013, Liu et al. proposed another type of complementary sequence sets, i.e. the quasi-complementary sequence sets (QCSSs), which could have larger set sizes than those of PCSSs [14]. In the literature, there are two kinds of QCSSs including periodic QCSSs and aperiodic QCSSs [14, 15]. In this paper, we mainly study periodic QCSSs. A periodic $(M,K,N,\vartheta_{\max})$-QCSS over an alphabet $A$ is a set consisting of $M$ two-dimensional $K\times N$ matrices such that the maximum value $\vartheta_{\max}$ of nontrivial periodic correlation sums of the row sequences is a small positive number. Each row of the two-dimensional matrices is called a constituent sequence of length $N$ and $K$ is called the flock size. The maximum non-trivial periodic correlation sum $\vartheta_{\max}$ measures the maximum possible amount of the multipath interference and the multiuser interference. For the application of a periodic QCSSs in MC-CDMA communications, it is required that its set size $M$ is as large as possible and $\vartheta_{\max}$ is as small as possible. However, there exists a tradeoff among the parameters $M,K,N$ and $\vartheta_{\max}$. In [14], Liu et al. presented a lower bound for the maximum periodic correlation sum of a periodic $(M,K,N,\vartheta_{\max})$-QCSS as follows:
$\vartheta_{\max}\geq\vartheta_{{\text{opt}}}=KN\sqrt{\frac{M/K-1}{MN-1}}.$ (1)
@@
It is easy to verify that $M\leq K$, which is the upper bound of the size of a periodic PCSS, provided that $\vartheta_{\max}=0$. The tightness factor of the above correlation lower bound for a periodic QCSS is defined by $\rho=\vartheta_{\max}/\vartheta_{{\text{opt}}}$. Note that $\rho\geq 1$. A QCSS is said to be optimal if $\rho=1$ and near-optimal if $1<\rho\leq 2$ [14]. Constructing optimal or near-optimal periodic QCSSs has been an interesting research topic in recent years. However, up to now, no infinite family of optimal QCSSs achieving the lower bound in (1) was reported. Only a few families of asymptotically optimal or near-optimal QCSSs have been constructed. In [14], Liu et al. constructed the first family of asymptotically optimal and the first family of asymptotically near-optimal periodic QCSSs from the Single difference sets and the existing optimal quaternary sequence sets. Using a similar idea, Li et al. constructed a family of asymptotically optimal periodic QCSSs from almost difference sets in [12]. In [9], Li et al. generalized the constructions in [14] and derived new asymptotically optimal periodic QCSSs. In [10, 11], Li et al. presented several constructions of asymptotically optimal QCSSs from additive and multiplicative characters over finite fields. Later, Luo et al. gave three new constructions of asymptotically optimal QCSSs with small alphabet sizes by additive characters over finite fields [16]. Recently, Xiao et al. used additive characters of finite fields to construct two more families of asymptotically optimal periodic QCSSs [22]. The parameters of these known families of periodic QCSSs are summarized in Table I. It is shown that only a small amount of known families of periodic QCSSs with set size $\Theta(K^{2})$ and flock size $K$ have been constructed and most of other known periodic QCSSs have set sizes much smaller than $\Theta(K^{2})$. The purpose of this paper is to use new approaches to construct five new families of asymptotically optimal or near optimal QCSSs with better parameters. The main contents of this paper are as follows:
  1. $\bullet$
@@
    Using a family of quadratic polynomials over finite fields, we construct a family of $(q^{2},q,q-1,q)$-QCSSs with alphabet size $p$ and a family of $(2^{2n},2^{n},2^{n}-1,2^{n})$-QCSSs with alphabet size $2$, where $q=p^{n}$ with $p$ an odd prime and $n$ a positive integer. These two families of QCSSs are asymptotically optimal with respect to the correlation lower bound and have larger set sizes than those in [10, 22] by Table I.
  2. $\bullet$
@@
    A family of $(2^{3n},2^{n},2^{n}-1,2^{n+1})$-QCSSs with alphabet size $2$ is constructed from a special quadrinomial over finite fields with characteristic 2. A family of $(p^{3n},p^{n},p^{n}-1,2p^{n})$-QCSSs with alphabet size $p$ is also constructed from a cubic quadrinomial over finite fields with characteristic $p$ for an odd prime $p$. These two families of QCSSs are asymptotically near-optimal with respect to the correlation lower bound. Note that the set size of each of them is the cube of the flock size. In this sense, these two families of QCSSs have the largest set sizes among all known families of asymptotically optimal or near-optimal periodic QCSSs.
  3. $\bullet$
@@
Table I: The parameters of known periodic QCSSs.
Set size $M$ Flock size $K$ Length $N$ $\vartheta_{\max}$ Alphabet Size Constraint References
@@
$p^{n}$ $\frac{p^{n}-1}{2}$ $p^{n}-1$ $\frac{p^{n}+1}{2}$ $p$ $p$ is an odd prime, $p^{n}>3$, [16]
$p^{n}$ $\frac{p^{n}-p^{n-1}}{2}$ $p^{n}-1$ $\frac{p^{n}+p^{n-1}}{2}$ $p$ $p$ is an odd prime, $p^{n}>3$, [16]
$p^{n}$ $u$ $p^{n}-1$ $v$ $p$ $p$ is an odd prime, $n>1$ is odd, [16]
$p^{2n}-p^{n}$ $p^{n}$ $p^{n}-1$ $p^{n}$ $p$ $p$ is a prime, $n>1$ [22]
$p^{2n}-p^{n}$ $p^{n}-1$ $p^{n}-1$ $p^{n}+1$ $p$ $p$ is a prime, $n>1$ [22]
@@
$p^{3n}$ $p^{n}$ $p^{n}-1$ $2p^{n}$ $p$ $p$ is an odd prime, $n\geq 1$, Theorem III.4
$p^{2n}-p^{n}$ $p^{n}-1$ $p^{n}-1$ $p^{n}$ $p(p^{n}-1)$ $p$ is a prime Theorem III.5
  • $u=\frac{p^{n}-p^{n-1}+(-1)^{(p-1)(n+3)/4}(p-1)p^{(n-1)/2}}{2}$, $v=\frac{p^{n}+p^{n-1}-(-1)^{(p-1)(n+3)/4}(p-1)p^{(n-1)/2}}{2}$,
    $w=(1+2^{n})\sqrt{2^{n}+2^{n-2}-2^{n/2}}$
@@ II Preliminaries
In this section, we recall some basic definitions of QCSSs and results on characters over finite fields, Gaussian sums, character sums, and the numbers of zeros of some polynomials over finite fields.
@@ II-A Periodic quasi-complementary sequence set
Let =$\mathbb{C}=\left\{\mathbf{C}^{0},\mathbf{C}^{1},\cdots,\mathbf{C}^{M-1}\right\}$ be a set of $M$ complementary sequences such that each $\mathbf{C}^{m}$ is a two-dimensional matrix with size $K\times N$ given by
@@
$R_{{{\textbf{C}}^{{m}_{1}}},{{\textbf{C}}^{{m}_{2}}}}(\tau)=\sum_{k=0}^{K-1}R_{{\mathbf{c}_{k}^{{m}_{1}}},{\mathbf{c}_{k}^{{m}_{2}}}}(\tau),\leavevmode\nobreak\ 0\leq\tau<N,$
@@
$\vartheta_{a}=\max\left\{|R_{{{\textbf{C}}^{{m}}},{{\textbf{C}}^{{m}}}}(\tau)|:0\leq m<M,\ 0<\tau<N\right\},$
$\vartheta_{c}=\max\left\{|R_{{{\textbf{C}}^{{m}_{1}}},{{\textbf{C}}^{{m}_{2}}}}(\tau)|:0\leq m_{1}\neq m_{2}<M,\ 0\leq\tau<N\right\}.$
@@
$\vartheta_{\max}=\max\{\vartheta_{c},\vartheta_{a}\}.$
It is an important performance measure of the complementary sequence set $\mathbb{C}$ in practice applications. If $\vartheta_{\max}>0$, then $\mathbb{C}$ is called a periodic $(M,K,N,\vartheta_{\max})$ quasi-complementary sequence set (QCSS). Otherwise, $\mathbb{C}$ is called a periodic perfect complementary sequence set (PCSS). Particularly, a PCSS reduces to a matrix consisting of two row sequences provided that $M=1$ and $K=2$. This matrix is called a Golay complementary pair.
@@ II-B Additive and multiplicative characters of finite fields
for each $j=0,1,\cdots,q-2$, defines a multiplicative character. In particular, $\varphi_{0}$ is called the trivial multiplicative character. For odd $q$, if $j=\frac{q-1}{2}$, then $\eta:=\varphi_{\frac{q-1}{2}}$ is called the quadratic multiplicative character of ${\mathbb{F}}_{q}$. The complex conjugate $\overline{\varphi}$ of a multiplicative character $\varphi$ is defined by $\overline{\varphi}(x)=\overline{\varphi(x)}=\varphi(x^{-1})$ for $x\in{\mathbb{F}}_{q}^{*}$. From [17], the orthogonal relation of multiplicative characters is given by
$\sum_{x\in{\mathbb{F}}_{q}^{*}}\varphi_{j}(x)=\left\{\begin{array}[]{ll}q-1&\mbox{if $j=0$},\\ 0&\mbox{if $1\leq j\leq q-2$}.\end{array}\right.$
@@ II-C Gaussian sums and character sums over finite fields
II-D The number of zeros of some polynomials over finite fields
In the following, we present some known results on the numbers of zeros of some polynomials over finite fields.
Lemma II.4.